# Hiding Functions within Functions: Steganography by Implicit Neural Representations

### v1

@@ VI-B2 Recoverability
We calculate the bit error rate (BER) between the original secret function ${{f}_{{\theta}}^{m}}$ and the secret function ${{f}_{{{\theta}_{e}}}^{m}}$ extracted from the stego function ${{f}_{\phi}^{c}}$. Table I presents the error rates of different architectures of secret function models in extracting secret models from the stego function. It can be seen that due to fixing the parameters of the secret function in the stego function, the secret function can be extracted without any loss, and the performance of the images represented by the secret function is not affected.
VI-B3 Fidelity
@@
Figure 8: The PSNR curves corresponding to the original carrier, the sampled image of the clean function, and the sampled image of the stego function, respectively.
Figure 9: The original carrier, the sampled image of the clean function, and the sampled image of the stego function.
We have used four network structures to evaluate the performance of the stego function, as shown in the second column of Table I. We also calculated the sampled results of the original cover image on both the clean model and the stego model, and used PSNR to measure their representation capabilities. The PSNR curves are shown in Fig 8, where PSNR12 represents the PSNR between the original image and the sampled image from the clean model, PSNR13 represents the PSNR between the original image and the sampled image from the stego function, and PSNR23 represents the PSNR between the sampled images from the clean model and the stego function, respectively. It can be observed that fixing some parameters indeed leads to a decrease in the representation capability of the stego function. As the parameters of the stego function model increase, the proportion of parameters of the secret function in the stego function decreases, resulting in an improvement in the representation capability of the stego function. Fig 9 presents the sampled results of the same cover image using different structures of clean functions and stego functions for representation. The first row represents the original cover image. The second row shows the sampled image obtained from the clean function, while the third row shows the sampled image obtained from the stego function.
Figure 10: PSNR curves under different expansion strategies.
Figure 11: Images Sampled from stego functions obtained by different expansion strategies.
We also evaluated the three different stego function strategies mentioned to verify the effectiveness of various function extension strategies. The three expansion strategies of the stego function are as follows: [64,64,64, 256,256,256],[256,256,256] and [256,256,256,256]. As shown in Fig 10, the stego function generated by the the hybrid mode exhibits the best performance. Fig11 show the sampled result under different modes. The first column represents the original cover image, and the second to fourth columns show the images obtained by sampling from the stego function constructed using horizontal, vertical, and hybrid modes, respectively.
VI-B4 Capacity
Our approach hides the parameters of one function within another function, and the capacity of the proposed method can be evaluated based on the expansion rate of the parameters. The function expansion rate is defined as:
@@ VI-B4 Capacity
where $N_{stego}$ and $N_{secret}$ represent the number of parameters in the stego function and secret function, respectively. Thus, we have $e\geqslant 1$. We further evaluate how the performance of our proposed method changes as the expansion rate varies. Fig 12 displays the implicit expressive ability of the stego function at different expansion rates.The structure of the secret function is [64, 64, 64], and the structure of the masking function is given in the second column of Table I.
Figure 12: The PSNR between the sampled images from the stego function and the original cover image under different expansion rates.
@@
VI-B5 Undetectability
In our method, the undetectability of the stego function (model) is a crucial concern. We establish a model pool that consists of multiple stego functions and their structurally identical clean functions. Both the stego and clean functions are trained to implicitly represent the same cover image. Since our method only requires training on a single image, we can create more clean and stego models. Ultimately, we construct a dataset consisting of 1200 stego functions and 1200 clean functions, each trained on 1200 different images.
Training a classifier to detect differences in parameter distributions between stego functions and clean functions is a direct approach. Following this idea, we randomly select 1000 parameters from each model as statistical measures of the parameters, resulting in a 1000-dimensional feature vector used to train a Support Vector Machine (SVM) classifier [30] . Specifically, we randomly select 1000 pairs of features from stego and clean functions for training, while the remaining 200 pairs are used for testing. The results in Table II indicate that the SVM method cannot detect the presence of secret functions from stego functions, with a detection accuracy close to $50\%$, where the SVM-poly refers to SVM with a polynomial kernel, while SVM-rbf refers to SVM with a Gaussian (RBF) kernel.
TABLE II: The Undetectability of the proposed method.
Detection method Detection rate
SVM-poly 0.53
SVM-rbf 0.55
Figure 13: The result on climate data.
VI-B6 Diversity of Data Representations
Implicit representation can be used to represent various types of data. We also conducted experiments on climate data from the ERA5 dataset [27]. The formats of ERA5 meteorological data are generally Grid and NetCDF formats. We convert Grib format data into npz data formats. In order to display the data intuitively, we save data as image format, as shown in the Fig 13. We use the data from January, April, July, October of 1979 as the secret message, and the temperature data from December of 2020 as the cover data. The structure of the secret function is [128, 128, 128], and the structure of the stego function is [512, 512, 512, 512]. We train all the models for 1000 epochs with $\eta=0.001$. From left to right, each column represents: original secret message data, secret message data sampled from the secret function, original cover data, sampled cover data from the stego function, and recovered secret message data from the stego function. From top to bottom, each row represents the experimental results using secret data from different months. The results show that with implicit representation, we can easily embed different types of message data into the stego function.
TABLE III: Comparasion with other Steganography
Category Method capacity:bpp/er BER/PSNR Eoob/Detection Rate PSNR/ACC
@@ VI-B6 Diversity of Data Representations
VI-C Comparison
As shown in Table III, we compared our proposed method with existing state-of-the-art multimedia (image) steganography schemes and model steganography schemes based on deep neural networks, evaluating them in terms of capacity, recoverability, detectability, and fidelity (only for model watermarking schemes). To ensure a fair comparison among different types of information hiding schemes, for multimedia steganography, we assume that the data to be hidden is a secret image of size $64\times 64$. For model steganography, we assume that the data is an MLP with a hidden layer size of [64,64,64].
For traditional steganography schemes [22][23][24] we copied the reported End-Of-Out-Band (EOOB) rate from each paper under a specific payload, where the payload refers to the number of bits per pixel, i.e., how many bits can be embedded per pixel. EOOB is calculated as the average of the false positive rate and the false negative rate.
For deep steganography schemes [59][57][58], we copied the results reported in Zhou et al.[5] for Perceptual Signal-to-Noise Ratio (PSNR) at 8bpp payload (on ImageNet) and detection accuracy.
@@ VI-C Comparison
VII Limitations and Future Work
Limitations: Although our simple approach can be regarded as a fusion of multimedia steganography and network model steganography, it does not outperform state-of-the-art image steganography or model steganography in terms of performance. In designing our method, we aimed for simplicity by directly utilizing function extensions to construct stego functions. The method of fixing the parameters of the secret function results in reduced performance when representing the cover data. Although increasing the expansion rate can improve the representational performance to some extent, the training time and the significant increase in parameters may raise concerns about this approach. Another fundamental issue is that, despite the considerable representational power of the implicit representation in theory, there is still a gap between implicit representation of images and other data and the original data. It may not be suitable for cases where lossless transmission of the secret message is required.

### v2

@@ 6.2.2 Recoverability
We calculate the bit error rate (BER) between the original secret function ${{f}_{{\theta}}^{m}}$ and the secret function ${{f}_{{{\theta}_{e}}}^{m}}$ extracted from the stego function ${{f}_{\phi}^{c}}$. Table. 1 presents the error rates of different architectures of secret function models in extracting secret models from the stego function. It can be seen that due to fixing the parameters of the secret function in the stego function, the secret function can be extracted without any loss, and the performance of the images represented by the secret function is not affected.
6.2.3 Fidelity
@@
Fig.  8: The PSNR curves corresponding to the original carrier, the sampled image of the clean function, and the sampled image of the stego function, respectively.
We have used four network structures to evaluate the performance of the stego function, as shown in the second column of Table 1. We also calculated the sampled results of the original cover image on both the clean model and the stego model, and used PSNR to measure their representation capabilities.
Fig.  9: The original carrier, the sampled image of the clean function, and the sampled image of the stego function.
The PSNR curves are shown in Fig. 8, where PSNR12 represents the PSNR between the original image and the sampled image from the clean model, PSNR13 represents the PSNR between the original image and the sampled image from the stego function, and PSNR23 represents the PSNR between the sampled images from the clean model and the stego function, respectively. It can be observed that fixing some parameters indeed leads to a decrease in the representation capability of the stego function. As the parameters of the stego function model increase, the proportion of parameters of the secret function in the stego function decreases, resulting in an improvement in the representation capability of the stego function. Fig. 9 presents the sampled results of the same cover image using different structures of clean functions and stego functions for representation. The first row represents the original cover image. The second row shows the sampled image obtained from the clean function, while the third row shows the sampled image obtained from the stego function.
Fig.  10: PSNR curves under different expansion strategies.
Fig.  11: The images sampled from stego functions obtained by different expansion strategies.
We also evaluated the three different stego function strategies mentioned to verify the effectiveness of various function extension strategies. The three expansion strategies of the stego function are as follows: [64,64,64, 256,256,256],[256,256,256] and [256,256,256,256]. As shown in Fig. 10, the stego function generated by the the hybrid mode exhibits the best performance. Figure 11 show the sampled result under different modes. The first column represents the original cover image, and the second to fourth columns show the images obtained by sampling from the stego function constructed using horizontal, vertical, and hybrid modes, respectively.
6.2.4 Capacity
Our approach hides the parameters of one function within another function, and the capacity of the proposed method can be evaluated based on the expansion rate of the parameters. The function expansion rate is defined as described in Eq. (16)
@@ 6.2.4 Capacity
where $N_{stego}$ and $N_{secret}$ represent the number of parameters in the stego function and secret function, respectively. Thus, we have $e\geqslant 1$. We further evaluate how the performance of our proposed method changes as the expansion rate varies. Fig 12 displays the implicit expressive ability of the stego function at different expansion rates.The structure of the secret function is [64, 64, 64], and the structure of the masking function is given in the second column of Table 1.
Fig.  12: The PSNR between the sampled images from the stego function and the original cover image under different expansion rates.
@@
6.2.5 Diversity of Data Representations
Implicit representation can be used to represent various types of data. We also conducted experiments on climate data from the ERA5 dataset[52]. The formats of ERA5 meteorological data are generally Grid and NetCDF formats. We convert Grib format data into npz data formats. In order to display the data intuitively, we save data as image format, as shown in the Fig. 13. We use the data from January, April, July, October of 1979 as the secret message, and the temperature data from December of 2020 as the cover data. The structure of the secret function is [128, 128, 128], and the structure of the stego function is [512, 512, 512, 512]. We train all the models for 1000 epochs with $\eta=0.001$. From left to right, each column represents: original secret message data, secret message data sampled from the secret function, original cover data, sampled cover data from the stego function, and recovered secret message data from the stego function. From top to bottom, each row represents the experimental results using secret data from different months. The results show that with implicit representation, we can easily embed different types of message data into the stego function.
On NeRF-Synthetic, we adopt [54] as the NeRF backbone architecture for efficiency. This network consists of four blocks, each block is composed of some linear layers and non-linear activation functions. The first stage of training for secret 3D message is performed according to the standard recipes of this implementation. We then extend the network that represents secret messages by maintaining four blocks and increasing the number of fully connected nodes in each block. The network that masks the carrier while keeping the original parameters unchanged is retrained. In order to display the 3D data, we save data as image format, as shown in the Fig. 14. Here we present two sets of experimental results, from top to bottom, representing stego 3D data using LEGO and representing secret 3D data using hot dogs, as well as stego 3D data using LEGO and representing secret data using 3D boat data. From the results, it can be observed that our method also performs effectively in representing implicit representations of 3D data. Similar to image representation, in the representation of disguised data, namely LEGO data, the rendering quality is somewhat inferior compared to the original LEGO data rendering quality.
Table 2: The Undetectability of the proposed method
Detection method Accuracy (Image) Accuracy (Climate Image) Accuracy (3D) Mean
SVM-poly 0.53 0.52 0.53 0.53
SVM-rbf 0.55 0.54 0.54 0.54
SVM-sigmoid 0.54 0.55 0.53 0.54
Table 3: Comparison with other Steganography Methods
Category Method Capacity (bpp/er) BER/PSNR Eoob/Detection Rate PSNR/ACC
@@ 6.2.5 Diversity of Data Representations
6.2.6 Undetectability
In our method, the undetectability of the stego function (model) is a crucial concern. We establish a model pool that consists of multiple stego functions and their structurally identical clean functions. Both the stego and clean functions are trained to implicitly represent the same cover (Image,Climate Image,3D data). Since our method only requires training on a single cover, we can create more nature and stego models. Ultimately, we construct a dataset consisting of 1200 stego functions and 1200 clean functions, each trained on 1200 different pairs of cover data,.
Training a classifier to detect differences in parameter distributions between stego functions and clean functions is a direct approach. Following this idea, we randomly select 1000 parameters from each model as statistical measures of the parameters, resulting in a 1000-dimensional feature vector used to train a Support Vector Machine (SVM) classifier [58]. Specifically, we randomly select 1000 pairs of features from stego and clean functions for training, while the remaining 200 pairs are used for testing. The results in Table 2 indicate that the SVM method cannot detect the presence of secret functions from stego functions, with a detection mean accuracy close to $50\%$, where the SVM-poly refers to SVM with a polynomial kernel, SVM-rbf refers to SVM with a Gaussian (RBF) kernel, while SVM-simoid refers to SVM with a sigmoid kernel.
Fig.  13: The result on climate data.
Fig.  14: The result on NeRF Synthetic 3D data.
6.3 Comparison
As shown in Table 3, we compared our proposed method with existing state-of-the-art multimedia (image) steganography schemes and model steganography schemes based on deep neural networks, evaluating them in terms of capacity, recoverability, detectability, and fidelity (only for model watermarking schemes). To ensure a fair comparison among different types of information hiding schemes, for multimedia steganography, we assume that the data to be hidden is a secret image of size $64\times 64$. For model steganography, we assume that the data is an MLP with a hidden layer size of [64,64,64].
For traditional steganography schemes[55, 20, 56] we copied the reported End-Of-Out-Band (EOOB) rate from each paper under a specific payload, where the payload refers to the number of bits per pixel (i.e., how many bits can be embedded per pixel). EOOB is calculated as the average of the false positive rate and the false negative rate.
For deep steganography schemes [57, 10, 12], we copied the results reported in Zhou[5] for Perceptual Signal-to-Noise Ratio (PSNR) at 8bpp payload (on ImageNet) and detection accuracy.
@@ 6.3 Comparison
7 Discussion
Limitations: While our method can be regarded as a fusion of multimedia steganography and network model steganography, it does not outperform state-of-the-art image or model steganography in terms of performance. In designing our method, we aimed for simplicity by directly utilizing function extensions to construct stego functions. However, the approach of fixing the parameters of the secret function results in reduced performance when representing the cover data. Although increasing the expansion rate can improve representational performance to some extent, the increased training time and significant rise in parameters may raise concerns about this approach.
Another fundamental issue is that, despite the considerable representational capability of implicit representation in theory, there remains a gap between the implicitly represented images and other data and the original data. However, due to the significant redundancy present in multimedia data like images, even if there is such a gap, steganography is generally considered successful as long as the semantic differences are small. However, in scenarios where the implicitly represented data must be completely identical to the original data, our approach is difficult to apply. For example, in the case of hiding text data, even if the implicit neural network overfits, it cannot completely and accurately represent text data due to its small redundancy and low tolerance for error. Even minor modifications can lead to semantic changes, making this method unsuitable for scenarios requiring lossless transmission of secret messages
@@ 8 Conclusion
Acknowledgment This work was supported in part by the National Key R&D Program of China under grant 2021YFB3100901, NSF of China under Grant 62074131, 62272478 and 62202496, and Shaanxi Provincial Key R&D Program 2023-ZDLGY-32.